OpenStudy (anonymous):

Solve -8x < 64. x < 8 x < -8 x > 8 x > -8

OpenStudy (anonymous):

64^(1/2) is the square root of 64 (i.e. 8) 64^(3/2) is the square root of 64 cubed 8^3 = 8*8*8 = 512 ANOTHER WAY to do this is to break 64 down into a product of prime factors: 64 = 8*8 ..... = 2*2*2*2*2*2 Hence 64^(3/2) = ((2^6)^(3/2)) Now the rule here is that to raise a power to a power, we multiply exponents. 6 * (3/2) = 9 Hence 64^(3/2) = 2^9 = 512

OpenStudy (anonymous):

remove the exponents n the roots!

OpenStudy (anonymous):

sorry for that

OpenStudy (anonymous):

....... lol

OpenStudy (anonymous):

sorry !

OpenStudy (kropot72):

x < -8